# UnlearnCanvas: A Stylized Image Dataset to Benchmark Machine Unlearning for Diffusion Models

### v2

@@ Abstract
The rapid advancement of diffusion models (DMs) has not only transformed various real-world industries but has also introduced negative societal concerns, including the generation of harmful content, copyright disputes, and the rise of stereotypes and biases. To mitigate these issues, machine unlearning (MU) has emerged as a potential solution, demonstrating its ability to remove undesired generative capabilities of DMs in various applications. However, by examining existing MU evaluation methods, we uncover several key challenges that can result in incomplete, inaccurate, or biased evaluations for MU in DMs. To address them, we enhance the evaluation metrics for MU, including the introduction of an often-overlooked retainability measurement for DMs post-unlearning. Additionally, we introduce UnlearnCanvas, a comprehensive high-resolution stylized image dataset that facilitates us to evaluate the unlearning of artistic painting styles in conjunction with associated image objects. We show that this dataset plays a pivotal role in establishing a standardized and automated evaluation framework for MU techniques on DMs, featuring 7 quantitative metrics to address various aspects of unlearning effectiveness. Through extensive experiments, we benchmark 5 state-of-the-art MU methods, revealing novel insights into their pros and cons, and the underlying unlearning mechanisms. Furthermore, we demonstrate the potential of UnlearnCanvas to benchmark other generative modeling tasks, such as style transfer. The UnlearnCanvas dataset, benchmark, and the codes to reproduce all the results in this work can be found at https://github.com/OPTML-Group/UnlearnCanvas.
Machine Learning
@@ 1 Introduction
In recent years, the field of text-to-image technology, epitomized by Diffusion Models (DMs), has seen remarkable progress, demonstrating an unprecedented ability to generate high-resolution, high-quality images across a wide array of subjects (Rombach et al., 2022; Midjourney et al.,, 2023). These models, widely accessible to public, find applications in diverse sectors, including advertising, the creative arts (Fotor et al.,, 2022), and forensic sketching (LabLab AI et al.,, 2023). Central to the training of DMs is their reliance on the data source from the Internet (Schuhmann et al., 2022) to cover a wide range of content. However, this inclusiveness comes with challenges, including potential concerns about harmful generation (Schramowski et al., 2023), copyright issues (Krietzberg, 2023), and biases/stereotypes (Vinker et al., 2023; Turk, 2023).
Figure 1: An illustration of the task MU for DMs. The pretrained model contains the generation capability of different concepts in different domains. MU aims to erase the generation of a certain unlearning target concept without harming the generation of others.
To alleviate the negative social impacts associated with DMs, machine unlearning (MU) techniques are catching increasing attention and have been studied in the field of vision generative models (Schramowski et al., 2023; Gandikota et al., 2023a, b; Kumari et al., 2023; Fan et al., 2023; Heng & Soh, 2023). MU for DMs (see Fig. 1 for an illustration) aims to erase the model’s capacity to generate images related to a specific concept, essentially erasing that concept from the pre-trained model. This concept can be provided in the form of text prompts (i.e., text conditions of DMs), covering a wide range from concrete object entities to abstract notions, such as artistic styles. We refer readers to Sec. 2 for more related work on MU for DMs.
Unlike MU studies on image classification (Izzo et al., 2021; Golatkar et al., 2020; Becker & Liebig, 2022; Thudi et al., 2022; Warnecke et al., 2021; Jia et al., 2023), we lack evaluation benchmarks for MU in DMs. The current primary choice for assessing the performance of unlearned DMs (i.e., DMs post unlearning) is the Inappropriate Image Prompts (I2P) dataset (Schramowski et al., 2023). This dataset consists of real-world image-to-text prompts designed for erasing concepts like ‘nudity’ and ‘violence’. However, I2P mainly contains isolated unlearning targets and focuses on safety issues related to image generation.
@@
To enhance the assessment of MU in DMs and establish a standardized evaluation framework, we propose to develop a new benchmark dataset, referred to as UnlearnCanvas, which is designed to evaluate the unlearning of artistic painting styles along with associated image objects. Unlike I2P, UnlearnCanvas can be used to assess the removal of copyrighted information, such as artistic style, and it allows for increased diversity by including various object classes and styles. We summarize our contributions below.
@@
$\bullet$ We conduct a systematic review of existing MU methods for DMs and identified three unresolved challenges in their evaluations, such as limited quantitative assessments and a lack of retainability evaluation of unlearned DMs.
$\bullet$ We propose UnlearnCanvas, a large-scale high-resolution stylized image dataset with 24,000 labeled images, 60+ styles, and 20+ object topics. Its style-object dual supervision and high stylistic consistency help us to improve the precision of MU evaluation for DMs.
$\bullet$ We develop a standardized evaluation pipeline for MU in DMs under UnlearnCanvas, which measures unlearning effectiveness, generation retainability, generation quality, and unlearning efficiency.
$\bullet$ We benchmark state-of-the-art MU methods for DMs using UnlearnCanvas, validating the significance of our proposed evaluation metrics and providing insights into the strengths, weaknesses, and unlearning mechanisms.
2 Related Work and Preliminaries
@@
Machine unlearning (MU) for generative models in vision. MU was originally developed to mitigate the (potentially detrimental) influence of specific data points in a pretrained ML model, without necessitating a complete retraining of the model after removing these unlearning data (Cao & Yang, 2015; Bourtoule et al., 2021; Xu et al., 2023). The significance of MU has emerged with the purpose of data privacy protection, in response to regulations such as ‘the right to be forgotten’ (Hoofnagle et al., 2019). However, it has rapidly gained recognition for its role in promoting security, safety, and trustworthiness of ML models. Examples include defense against backdoor attacks (Liu et al., 2022; Jia et al., 2023), fairness enhancement (Chen et al., 2023; Oesterling et al., 2023), and controllable federated learning (Halimi et al., 2022; Jin et al., 2023).
More recently, unlearning has proven crucial in text-to-image generation via DMs to prevent the production of harmful, private, or illegal content (Gandikota et al., 2023a; Zhang et al., 2023a; Gandikota et al., 2023b; Fan et al., 2023; Kumari et al., 2023; Heng & Soh, 2023). For example, DMs have demonstrated a susceptibility to generating NSFW (Not Safe For Work) images when conditioned on inappropriate text prompts (e.g., ‘nudity’ and ‘violence’) (Schramowski et al., 2023; Zhang et al., 2023b; Chin et al., 2023). Similarly, MU has also found applications in preventing the generation of copyrighted information, such as the misuse of artistic painting styles (Gandikota et al., 2023a; Zhang et al., 2023a). In summary, MU can be regarded as a (post-training) model edit operation, aimed at preventing DMs from producing undesired or inappropriate images.
@@
Our setup and objective. Let us denote a pretrained DM by $\bm{\theta}_{\mathrm{o}}$, capable of synthesizing high-quality images conditioned on a provided text prompt $c$ (e.g., ‘A painting of a cat in Van Gogh style’). In this work, we focus on the task of generating images in various painting styles. When there is a request to adapt the model $\bm{\theta}_{\mathrm{o}}$ to prevent the generation of images, e.g., in the ‘Van Gogh’ style, the MU problem arises, with the ‘Van Gogh’ style being the specified unlearning target, also known as the erasing concept (Gandikota et al., 2023a). See Fig. 1 for a visual representation of the concept.
In formulating MU methods, the traditional setup presupposes access to the training dataset $\mathcal{D}$, where $\mathcal{D}_{\mathrm{f}}\subset\mathcal{D}$ represents the forget set that revolves around the unlearning target. A commonly used gold standard of MU is to retrain the model from scratch over the retain set $\mathcal{D}_{\mathrm{r}}=\mathcal{D}\setminus\mathcal{D}_{\mathrm{f}}$. However, commencing the training process from scratch becomes practically unfeasible because of (1) the substantial computational expense it entails, (2) the difficulty in accurately identifying $\mathcal{D}_{\mathrm{f}}$, and (3) the absence of access to the original training data source.
To overcome the aforementioned difficulties (1)-(3), the current research mainstream (Gandikota et al., 2023a; Kumari et al., 2023; Zhang et al., 2023a; Gandikota et al., 2023b; Heng & Soh, 2023; Fan et al., 2023) has focused on developing approximate yet efficient unlearning methods as practical alternatives to full retraining. The approximation introduced by current unlearning methods raises the question that we aim to address in this work:
How can we accurately assess unlearning performance and establish a benchmark for existing unlearning methods?
3 Unresolved Challenges in MU Evaluation
This work will develop a new benchmark dataset, referred to as UnlearnCanvas (detailed in Sec. 4), focusing on the task of unlearning artistic painting styles in conjunction with associated image objects. With the use of UnlearnCanvas, a text prompt for DMs can be constructed as ‘A painting of a {object} in {style}’. And the unlearning target can be explicitly designated as either ‘style’ or ‘object’. To be specific, UnlearnCanvas is targeted for the open challenges (C1-C3) when building a rigorous quantitative evaluation framework for MU in DMs. These challenges are consolidated through a careful examination of existing studies for object and/or style unlearning in Tab. 1.
Table 1: An overview of the scope and evaluation methods of the existing MU methods for DMs. The symbol denotes neither objects nor styles are covered.
Related work Unlearning targets Quantitative evaluation
Unlearning
effectiveness
Retainability
ESD (Gandikota et al., 2023a)
5 artistic styles
11 objects
objects only objects only
CA (Kumari et al., 2023)
4 artistic styles
4 objects
objects only
FMN (Zhang et al., 2023a)
1 artistic style
7 objects
objects only
UCE (Gandikota et al., 2023b)
7 artistic settings [^1] UCE adopted a multi-concept unlearning setting, where different numbers of styles are unlearned at the same time and total up to 1000 styles are considered.
11 objects
objects only objects only
SA (Heng & Soh, 2023)
1 artistic style
31 objects
objects only
SalUn (Fan et al., 2023)
No artistic styles
11 objects
objects only objects only
(C1) The absence of a diverse yet cohesive unlearning target repository. As shown in Tab. 1, the assessment of MU for DMs, in terms of both effectiveness (concept erasure performance) and retainability (preserved generation quality under non-forgotten concepts), is typically conducted using manually selected targets, often chosen from a limited pool of unlearning targets. A critical aspect missing from MU evaluation is the recognition that unlearning targets related to styles and objects should not be isolated within the dataset used for MU assessment. Instead, they should be considered in a manner that exposes their potential interconnected influences on other targets.
Table 2: Classification accuracy of a style classifier (ViT-L/16) finetuned on WikiArt (Phillips & Mackintosh, 2011). The classifier is tested on the test set and the image set generated by SD v1.5 and v2.0 with the prompt ‘An painting in {artist} style’, where the {artist} is traversed over the WikiArt labels (10 images/prompt).
WikiArt
Train Set
WikiArt
Test Set
Images by
SD v1.5
Images by
SD v2.0
Accuracy 99.8% 85.4% 56.7% 41.2%
(C2) The lack of precision in evaluation. Artistic styles can be challenging to precisely define and distinguish, which makes quantifying their impact on unlearning effectiveness and retainability difficult. To illustrate this challenge, we conducted a preliminary investigation in Tab. 2 using a ViT-based classifier trained on the WikiArt dataset (Phillips & Mackintosh, 2011). This dataset categorizes paintings into 139 distinct styles, each annotated with the creator’s name. Despite training, the ViT-large model achieved only an $84.32\%$ accuracy on the test dataset. This result underscores the inherent complexity of classifying artistic styles, causing inaccuracies in quantitative MU evaluations. The classifier’s suboptimal accuracy on DMs generated images further increases the prediction error (around $50\%$ accuracy). This complicates precise quantitative evaluations.
(C3) The lack of a systematic study on ‘retainability’ of DMs post unlearning. As indicated by the ‘retainability’ column in Tab. 1, assessing the capacity of unlearned DMs to maintain image generation under innocent concepts is notably lacking. However, retainability is a critical indicator of potential ‘side effects’ of an unlearning method. In conjunction with challenge (C1), the inclusion of diverse unlearning targets, encompassing both styles and objects, allows for the evaluation of retainability from various perspectives. For instance, when unlearning an artistic style, it is crucial to consider the retainability of image generation abilities for other styles as well as for objects.
The challenges (C1)-(C3) outlined above have driven us to develop UnlearnCanvas as a solution for the systematic and comprehensive evaluation of MU in DMs. We will show in Sec. 4 that UnlearnCanvas includes a diverse range of artistic styles and objects. Within this dataset, we can explore the interrelationships between paired image styles and image objects to address (C1). Each style category in UnlearnCanvas maintains high intra-style consistency, facilitating accurate style and object classification and enabling post-unlearning quantitative evaluation for (C2). To address (C3), we can evaluate two types of retainability, providing a precise and comprehensive assessment of the potential side effects induced by MU methods.
@@ 4 Our Proposal: UnlearnCanvas Dataset
Construction of UnlearnCanvas. As motivated in Sec. 2, UnlearnCanvas is created for ease of MU evaluation in DMs. Its construction involves two main steps: seed image collection and subsequent image stylization; see Fig. 2 for a schematic overview.
@@
Seed image collection. UnlearnCanvas includes images in $60$ unique artistic styles across $20$ distinct objects. The style categories are derived from a set of high-resolution photos, termed seed images. These seed images consist of original real-world photos from Pexels (Pexels et al.,, 2023). In total, there are 400 seed images across 20 object classes.
@@
Image stylization. After collecting the seed images, the next step in dataset development is the stylization process. During this phase, each seed image is transformed into 60 predetermined artistic styles. The goal is to ensure that the resulting images maintain high stylistic consistency while preserving the content details present in the seed images. This stylization process is carried out using the services provided by (Fotor et al.,, 2022). Once all the seed images have been stylized in all 60 styles, the dataset is constructed with super-high-resolution images and organized in a hierarchical structure that balances both style and object categories.
Figure 2: An illustration of the key steps when curating UnlearnCanvas and its key features.
@@
Advantages of UnlearnCanvas. Next, we will highlight the advantages (A1-A3) of UnlearnCanvas to address the MU evaluation challenges (C1-C3) discussed in Sec. 3.
@@
(A1) Style-object dual supervision enables a rich unlearning target bank for comprehensive evaluation. UnlearnCanvas offers a diverse range of unlearning targets by annotating each image with both style and object labels. This is in response to (C1). We have categorized the unlearning targets into two main domains: a style domain comprising $60$ artistic styles, and an object domain including $20$ object classes. This setup allows us to investigate the impact of style unlearning on the fidelity of generated objects in images and vice versa. Besides its role in MU evaluation, UnlearnCanvas could also serve as an annotated testbed for assessing style transfer accuracy and visual in-context learning performance; see details in Appx. C.
(A2) High stylistic consistency ensures precise style definitions and enables accurate quantitative evaluations. In UnlearnCanvas, images labeled with the same style exhibit high intra-style consistency and distinct inter-style differences, as exemplified in Fig. 2. This facilitates the distinction between different styles, effectively addressing challenge (C2). In a similar study to that presented in Tab. 2, we observe that the styles within UnlearnCanvas can be readily classified using ViT-Large (Dosovitskiy et al., 2020), achieving an accuracy of $98\%$ for images generated by SD finetuned on UnlearnCanvas. We refer readers to Tab. A2 for detailed results.
Figure 3: Evaluation metric illustration, with unlearning the ‘Van Gogh style’ as a case study. ✓ and ✗ denote desired and undesired results post unlearning.
@@
(A3) Enabling in-depth retainability analyses for MU evaluation. Using UnlearnCanvas, we can assess DMs’ retainability post-unlearning from two perspectives: in-domain retainability, which evaluates the model’s performance on test-time generations within the same domain as the unlearning target, and cross-domain retainability, which assesses the model’s ability to generate concepts from a different domain. For example, as depicted in Fig. 3, we can define the ‘domain’ as either ‘artistic style’ or ‘object class’. If the unlearning target is the ‘Van Gogh’ style (i.e., the concept to be erased), then generating images in all styles except ‘Van Gogh’ but within the same object class (i.e., ‘Dog’ in Fig. 3) represents the in-domain case. Conversely, out-of-domain refers to generating images with different object classes (e.g., ‘Horses’ in Fig. 3), i.e., objects different from the one associated with the unlearning target.
Figure 4: Image examples with the label “Van Gogh Style” from WikiArt (Phillips & Mackintosh, 2011) and UnlearnCanvas. Images of the same artistic style in UnlearnCanvas exhibit high stylistic consistency compared to WikiArt.
Figure 5: An illustration of the evaluation pipeline proposed in this work using UnlearnCanvas when unlearning a specific target concept ‘Van Gogh Style’. Unlearning performances (including the unlearning effectiveness and retainability) are quantitatively assessed (marked in blue) to accurately reflect the unlearning performance portrait of each method. The unlearning target of the pipeline could traverse all the styles and objects to achieve a complete and comprehensive evaluation.
UnlearnCanvas vs. WikiArt.
To the best of our knowledge, WikiArt (Nichol, 2016) is the most relevant baseline dataset to ours. In Tab. A1, we provide a direct comparison of the key attributes of these two datasets. UnlearnCanvas differs from WikiArt in the following aspects. First, UnlearnCanvas includes a greater number of high-resolution images ($15$M) compared to WikiArt ($2$M), a factor that may enhance the training of state-of-the-art DMs. Second, UnlearnCanvas surpasses WikiArt in terms of both intra-style coherence and inter-style distinctiveness, as illustrated in Fig. 4, where images labeled with ‘Van Gogh Style’ from both datasets are compared. In UnlearnCanvas, the images exhibit high stylistic consistency, while WikiArt lacks the necessary clarity for precise assessment. This will hamper the MU evaluation as discussed in the challenges (C2) and (C3).
5 Evaluation Pipeline
@@
In this section, we introduce the evaluation pipeline and the benchmarked MU methods for UnlearnCanvas. Fig. 5 illustrates the proposed evaluation pipeline for MU in DMs with a designated unlearning target. It comprises four phases (I-IV) to evaluate unlearning effectiveness, retainability, generation quality, and efficiency.
@@
Phase I: Testbed preparation (Fig. 5a). Prior to unlearning, we commence by fine-tuning a specific DM, given by SD (StableDiffusion) v1.5, on UnlearnCanvas for text-to-image generation. We also train a vision transformer (ViT-Large) (Dosovitskiy et al., 2020) for the purposes of style and object classification after unlearning. We ensure that both the SD model and classifiers attain high performance in image generation and classification across all styles and objects. See detailed testbed preparation in Appx. A.

### v4

# UnlearnCanvas: A Stylized Image Dataset for Enhanced Machine Unlearning Evaluation in Diffusion Models

@@ Abstract
The technological advancements in diffusion models (DMs) have demonstrated unprecedented capabilities in text-to-image generation and are widely used in diverse applications. However, they have also raised significant societal concerns, such as the generation of harmful content and copyright disputes. Machine unlearning (MU) has emerged as a promising solution, capable of removing undesired generative capabilities from DMs. However, existing MU evaluation systems present several key challenges that can result in incomplete and inaccurate assessments. To address these issues, we propose UnlearnCanvas, a comprehensive high-resolution stylized image dataset that facilitates the evaluation of the unlearning of artistic styles and associated objects. This dataset enables the establishment of a standardized, automated evaluation framework with 7 quantitative metrics assessing various aspects of the unlearning performance for DMs. Through extensive experiments, we benchmark 9 state-of-the-art MU methods for DMs, revealing novel insights into their strengths, weaknesses, and underlying mechanisms. Additionally, we explore challenging unlearning scenarios for DMs to evaluate worst-case performance against adversarial prompts, the unlearning of finer-scale concepts, and sequential unlearning. We hope that this study can pave the way for developing more effective, accurate, and robust DM unlearning methods, ensuring safer and more ethical applications of DMs in the future. The dataset, benchmark, and codes are publicly available at https://unlearn-canvas.netlify.app/.
@@ 1 Introduction
The recent technological breakthroughs in text-to-image generation, driven by diffusion models (DMs), have shown an unprecedented capability to produce high-resolution, high-quality images across a diverse range of subjects [1, 2, 3, 4, 5, 6, 7, 8, 9, 10]. These models have become widely accessible to the public and are applied across various sectors, including advertising, the creative arts [11], and forensic sketching [12]. One reason for DMs being able to generate a broad spectrum of content is their reliance on diverse internet-sourced data [13]. However, this inclusiveness comes with risks, such as harmful generation [14], copyright issues [15], and biases or stereotypes [16, 17].
Refer to caption    Refer to caption
(a) (b)
Figure 1: (a) An illustration of MU for DMs. (b) Overview of experiment settings and benchmark results. This benchmark focuses on three categories of quantitative metrics: the unlearning effectiveness (UA, Rob., FU, SU); the retainability of innocent knowledge (IRA, CRA, FR, SR); and the image generation quality (FID). Results are normalized to $0\%\sim 100\%$ per metric. No single method excels across all metrics. See a summary of these metrics in Tab. A1 and more results in Sec. 4.
To alleviate the negative social impacts associated with DMs, machine unlearning (MU) techniques are catching increasing attention and have been studied in the field of text-to-image generation via DMs [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33]. MU for DMs, also referred to as DM unlearning [27], aims to prevent the model from generating images when conditioned on an undesired concept (typically specified by a text prompt to be forgotten). Therefore, the problem of DM unlearning is also known as concept erasing [18, 19, 20, 21, 22, 23, 24, 25, 26]. Based on the types of unlearning requests, the targeted concept to be erased can be diverse, including not-safe-for-work (NSFW) prompts [14, 23], concrete object entities [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33], and copyrighted information like artistic styles [23, 24, 25, 26, 28, 29, 30, 32]. In Fig. 1 (a), we provide an illustration of DM unlearning by demonstrating the removal of an undesired artistic style from DM-generated images. Additionally, we refer readers to Sec. 2 for more related work on DM unlearning.
Figure 2: An illustration of machine unlearning using UnlearnCanvas. Concepts in the knowledge bank are categorized into different domains (style and object) and serve as potential unlearning targets. When one concept is unlearned, the rest concepts in both the same and different domains are required to be retained.
Compared to MU for DMs, the unlearning studies in computer vision have primarily focused on image classification models [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. Given an image classification dataset, benchmarking unlearning performance for discriminative models is simpler, because MU typically involves either class-wise forgetting [49, 35] or data-wise forgetting [27, 40]. The former aims to eliminate the influence of a specific image class, while the latter removes the influence of specific data points from the entire training set. In contrast, DM unlearning is often applied to scenarios involving the removal of higher-level and more abstract concepts, rather than specific training data points, which can also be difficult to localize within the DM training set. The primary method for assessing the performance of DM unlearning is using the I2P (inappropriate image prompts) dataset [14]. This dataset mainly focuses on the safety assessment of unlearned DMs (i.e., DMs post-unlearning), designed to erase specific harmful concepts like ‘nudity’ and ‘violence’. Although I2P specifies which concepts to unlearn, evaluating the unlearning effectiveness and generation retainability of unlearned DMs under I2P remains challenging. This difficulty arises primarily from the lack of ground-truth data or objective criteria to precisely define safety versus non-safety. Furthermore, I2P cannot evaluate the performance of DM unlearning in other scenarios, such as object unlearning [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and style unlearning [23, 24, 25, 26, 27, 28, 29, 30].
@@
To enhance the assessment of machine unlearning in DMs and establish a standardized evaluation framework, we propose the development of a new benchmark dataset, referred to as UnlearnCanvas. Unlike I2P, UnlearnCanvas is designed to evaluate the unlearning of artistic painting styles along with associated image objects (see an illustration in Fig. 2). It also provides ground-truth data annotations to precisely define the criteria for the effectiveness of unlearning and the retained model utility post-unlearning (see a result overview in Fig. 1 (b)). We summarize our contributions below.
@@
$\bullet$ We conduct a systematic review of existing MU methods for DMs and identified three unresolved challenges in their evaluations. We further provide an in-depth analysis of why a standardized benchmark for DM unlearning evaluation is crucial.
$\bullet$ We propose UnlearnCanvas, a large-scale, high-resolution stylized image dataset with diverse styles and objects, to address current challenges in DM unlearning evaluation. Its dual supervision of styles and objects ensures stylistic consistency, enabling a comprehensive, standardized evaluation approach to precisely characterize unlearning efficacy, generation utility preservation, and efficiency.
$\bullet$ We benchmark 9 state-of-the-art DM unlearning methods using UnlearnCanvas, covering standard assessments and examining more challenging scenarios, such as adversarial robustness, grouped object-style unlearning, and sequential unlearning. This comprehensive evaluation provides previously unknown insights into their strengths and weaknesses.
2 Related Work, Problem Statement, and Open Challenges
Related work.
@@ 2 Related Work, Problem Statement, and Open Challenges
MU (machine unlearning) was originally developed to mitigate the (potentially detrimental) influence of specific data points in a pretrained ML model, without necessitating a complete retraining of the model after removing these unlearning data [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. The significance of MU has emerged with the purpose of data privacy protection, in response to regulations such as ‘the right to be forgotten’ [53]. However, it has rapidly gained recognition for its role in promoting security, safety, and trustworthiness of ML models. Examples include defense against backdoor attacks [38, 54], fairness enhancement [55, 56], and controllable federated learning [57, 58]. More recently, DM unlearning has proven crucial in text-to-image generation to prevent the production of harmful, private, or illegal content [23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. For example, DMs have demonstrated a susceptibility to generating NSFW (not-safe-for-work) images when conditioned on inappropriate text prompts (e.g., ‘nudity’ and ‘violence’) [14, 59, 60]. Similarly, MU has also found applications in preventing the generation of copyrighted information, such as the misuse of artistic painting styles [23, 28]. In brief, DM unlearning can be regarded as a model edit operation, aimed at preventing DMs from producing undesired or inappropriate images.
As unlearning approaches continue to evolve, the need for benchmarking their performance becomes increasingly critical. Standard datasets like CIFAR-10 [61] and ImageNet [62] have long been used to evaluate unlearning performance when forgetting a specific subset of data points [27, 40] or a targeted image class [49, 35]. Additionally, a recent benchmark specialized for unlearning in face recognition has been developed [63]. However, crafting effective benchmarks for generative models introduces additional challenges. In the language domain, recent initiatives have specifically aimed to evaluate the efficacy of MU for large language models (LLMs). Notable examples include TOFU [64], which explores the unlearning of fictitious elements, and WMDP [65], aimed at removing hazardous knowledge from LLMs. In contrast, to the best of our knowledge, there is currently no effective benchmark dataset designed for DM unlearning in text-to-image generation. The predominant dataset, I2P [14], targets the removal of harmful concepts for safe image generation. However, I2P only covers a specialized unlearning scenario and lacks precise evaluation criteria to support a comprehensive, quantitative, and precise assessment of unlearning performance. This gap inspires us to develop a new benchmark dataset for DM unlearning.
Table 1: Overview of existing MU evaluations for DMs, which covers two unlearning scenarios: object unlearning and style unlearning. Each case is demonstrated by the number of targeted unlearning concepts (target #), qualitative evaluation (Qual.) via generated image visualization, and quantitative evaluation (Quant.) of unlearning effectiveness (UE) and retainability (RT).
Related Work Object Unlearning Style Unlearning
Target # Quant. Qual. Target # Quant. Qual.
UE RT UE RT
ESD [23] 11 5
CA [16] 4 4
FMN [28] 7 1
UCE [24] 11 7
SA [29] 31 1
SalUn [27] 11 0
SPM [26] 24 5
SEOT [30] 22 2
EDiff [31] 2 0
SHS [32] 1 0
@@
Problem statement. Let us denote a pretrained DM by $\bm{\theta}_{\mathrm{o}}$, capable of synthesizing high-quality images conditioned on a provided text prompt $c$ (e.g., ‘A painting of a cat in Van Gogh style’). When there is a request to adapt the model $\bm{\theta}_{\mathrm{o}}$ to prevent generating images in a specific concept, such as the ‘Van Gogh’ painting style, the MU problem arises. Here, the ‘Van Gogh’ style becomes the specified unlearning target, also known as the erasing concept [23]. See Fig. 1-(a) for a visual representation of this unlearning problem. In this work, we specify the unlearning target as artistic styles and/or object classes to encompass the settings of style unlearning and object unlearning as described in the literature [23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
Motivation: Unresolved challenges in MU evaluation. When building a rigorous quantitative evaluation framework for DM unlearning, we identified three urgent challenges (C1)-(C3) after carefully examining existing studies as shown in Tab. 1.
(C1) The absence of a consensus on a diverse unlearning target test repository. As shown in Tab. 1, assessments of DM unlearning in various studies, in terms of both effectiveness (i.e., concept erasure performance) and retainability (i.e., preserved generation quality under non-forgotten concepts), typically use manually-selected unlearning targets from a limited pool and focus on object-centric evaluation. Precise evaluation for style unlearning and grouped object-style unlearning is lacking. Although the existing dataset WikiArt [66] features a collection of real-world artworks by various artists, it does not apply to object unlearning.
(C2) The lack of a systematic study on ‘retainability’ of DMs post-unlearning. As highlighted in the ‘retainability’ column of Tab. 1, there is a notable deficiency in the quantitative assessment of retainability for DMs post-unlearning. Retainability is crucial for measuring the potential side effects of a MU method. For instance, when unlearning an artistic style, it is essential to assess the DM’s ability to generate images in other styles and all objects. This makes the WikiArt dataset less suitable for benchmarking DM unlearning, as it lacks object labels needed to characterize the side effects of style unlearning on object recognition.
(C3) The precision challenge in evaluating DM-generated images. Artistic styles are inherently complex to define and differentiate precisely, complicating the quantitative evaluation of forgetting and retaining performance in style unlearning. For example, the WikiArt dataset was recently used in [67] to train a classifier for detecting copyright infringement, merely achieving an accuracy of 72.80%. The difficulty in precisely recognizing artistic styles in WikiArt may further hamper the evaluation of DMs’ post-unlearning generation. Tab. A3 illustrates the challenge of accurately recognizing the styles of images generated by DMs, even when finetuned on WikiArt. This is evidenced by the test-time classification accuracy using the style classifier, ViT-L/16 [68], finetuned on WikiArt. As we can see, only about half of the images are correctly classified into their corresponding style classes, e.g., 56.7% accuracy for classification on generated images using finetuned stable diffusion (SD) v2.0 [1], which is significantly lower than that on test images from the original WikiArt (85.4%).
The above challenges (C1)-(C3) drive us to develop the UnlearnCanvas dataset as a solution for the systematic and comprehensive evaluation of DM unlearning in Sec. 3.
@@ 3 Our Proposal: UnlearnCanvas Dataset
Construction of UnlearnCanvas. As motivated in Sec. 2, UnlearnCanvas is created for ease of MU evaluation in DMs. Its construction involves two main steps: seed image collection and subsequent image stylization; see Fig. 3 for a schematic overview. More information about the dataset is provided in Appx. A.
Figure 3: Illustration of curating UnlearnCanvas.
@@
Seed image collection. UnlearnCanvas includes images in $60$ unique artistic styles across $20$ distinct objects. The style categories are built upon a set of high-resolution seed images, given by real-world photos from Pexels [69]. There are $20$ seed images collected for each of the $20$ object classes. See step 1 in Fig. 3.
@@
Image stylization. After collecting the seed images, the next step is the stylization process. During this phase, each seed image is transformed into 60 predetermined artistic styles, ensuring high stylistic consistency while preserving the original content details. This stylization process is carried out using services provided by Fotor [11]. Once all seed images have been stylized in all 60 styles, the dataset is constructed with super-high-resolution images and organized in a hierarchical structure that balances both style and object categories. See step 2 in Fig. 3.
@@
Advantages of UnlearnCanvas. Next, we will elaborate on the advantages (A1)-(A3) of UnlearnCanvas to address the MU evaluation challenges (C1)-(C3) discussed in Sec. 2.
@@
(A1) Style-object dual supervision enables a rich unlearning target bank for comprehensive evaluation. To address (C1), UnlearnCanvas offers a diverse range of unlearning targets by annotating each image with both style and object labels. We categorized the unlearning targets into two main domains: a style domain comprising 60 styles, and an object domain with 20 object classes. This setup allows us to investigate the impact of style unlearning on the fidelity of object generation and vice versa.
@@
(A2) Enabling both ‘in-domain’ and ‘cross-domain’ retainability analyses for MU evaluation. To address (C2), UnlearnCanvas facilitates us to assess DMs’ retainability post-unlearning from two perspectives: in-domain retainability, which evaluates the DM’s generation performance within the same domain as the unlearning target, and cross-domain retainability, which assesses the model’s generation ability under conditions from a different domain. For example, as depicted in Fig. 4, we can define the ‘domain’ as either ‘artistic style’ or ‘object class’. If the unlearning target is the ‘Van Gogh’ style, then generating images in all styles except ‘Van Gogh’ within the same object class (e.g., ‘Dog’ in Fig. 4) represents the in-domain case. Conversely, cross-domain retainability refers to generating images of different object classes (e.g., ‘Horses’ in Fig. 4), corresponding to an object domain different from the unlearning target within the style domain.
Figure 4: Illustration of in-domain and cross-domain retainability evaluation, with the Van Gogh style as the unlearning target. ✓ and ✗ indicate satisfactory and undesired results post unlearning.
(A3) High stylistic consistency ensures precise style definitions and enables accurate quantitative evaluations. In UnlearnCanvas, images labeled with the same style exhibit high intra-style consistency and distinct inter-style differences. This facilitates the distinction between different styles, effectively addressing challenge (C2). In a similar study to that presented in Tab. A3, we observe that the styles within UnlearnCanvas can be readily classified using ViT-Large [68], achieving an accuracy of 98% for images generated by DM finetuned on UnlearnCanvas. For detailed results, please refer to Tab. A4.
The above advantages (A1)-(A3) make UnlearnCanvas uniquely suitable for assessing the performance of DM unlearning. For instance, UnlearnCanvas includes a greater number of high-resolution images (15M) compared to the existing dataset WikiArt (2M). Moreover, UnlearnCanvas surpasses WikiArt in terms of both intra-style coherence and inter-style distinctiveness. See Appx. C for more detailed comparisons.
@@
Evaluation pipeline via UnlearnCanvas. We next introduce how UnlearnCanvas can be used to benchmark the performance of DM unlearning and the corresponding evaluation metrics. Fig. 5 illustrates the proposed evaluation pipeline with a designated unlearning target. It comprises four phases (I-IV) to evaluate unlearning effectiveness, retainability, generation quality, and efficiency.
@@
Phase I: Testbed preparation (Fig. 5 (a)). Prior to unlearning, we commence by finetuning a specific DM, given by SD (StableDiffusion) v1.5, on UnlearnCanvas for text-to-image generation. We also train a vision transformer (ViT-Large) [68] for the purposes of style and object classification after unlearning. We ensure that both the SD model and classifiers attain high performance in image generation and classification across all styles and objects. See detailed testbed preparation in Appx. B.